# Ram-EHR: Retrieval Augmentation Meets Clinical Predictions on Electronic Health Records

### v2

@@ Abstract
Ran Xu1thanks:  Equal contribution., Wenqi Shi211footnotemark: 1, Yue Yu2, Yuchen Zhuang2, Bowen Jin3 May D. Wang2, Joyce C. Ho1, Carl Yang1 1 Emory University  2 Georgia Institute of Technology 3 University of Illinois at Urbana Champaign {ran.xu,joyce.c.ho,j.carlyang}@emory.edu, {wshi83,yczhuang,yueyu,maywang}@gatech.edu, bowenj4@illinois.edu
@@ 1 Introduction
To further improve the downstream predictive performance, several works attempt to augment the EHR visits with external knowledge. For example, van Aken et al. (2021) and Naik et al. (2022) incorporate additional clinical notes, although these clinical notes can be noisy and contain irrelevant contents for clinical predictions; another solution is to leverage external clinical knowledge graphs (KGs), such as UMLS (Chandak et al., 2023), which contains rich medical concepts (e.g., disease, medications) and their corresponding relationships (Jin et al., 2023). Integrating KGs with EHRs has been shown to boost model performance (Choi et al., 2017; Xu et al., 2023b; Gao et al., 2023). However, these works mostly rely on knowledge from a single source and medical KGs mainly focus on specific types of relations (e.g., hierarchical relations), which do not comprehensively capture the semantic information for medical codes (e.g., phenotype). Besides, it is non-trivial to align medical codes in EHRs with KGs due to the non-uniformity of surface names (e.g., abbreviations or colloquial terms) (Hao et al., 2021; Zhang et al., 2022). Jiang et al. (2024) use knowledge generated from large language models (LLMs) to assist EHR prediction, but LLMs may not always provide the most relevant knowledge for target tasks and face the risk of hallucination. Effectively leveraging external knowledge to facilitate EHR predictive tasks remains a significant challenge.
@@ 2 Related Works
Retrieval augmented learning, which collects additional contextual information from external corpus, has shown effectiveness on diverse tasks including language modeling (Borgeaud et al., 2022; Wang et al., 2023a), knowledge-intensive NLP (Lewis et al., 2020; Shi et al., 2023a), commonsense reasoning (Yu et al., 2022a), code genereation (Parvez et al., 2021), and few/zero-shot learning (Xu et al., 2023a; Izacard et al., 2023). Compared to the general domain, the application of retrieval-augmented learning to clinical tasks is still under-explored. Some efforts have been paid to retrieval augmented clinical language models (Zakka et al., 2024), clinical decision making (Shi et al., 2023b; Thompson et al., 2023) and clinical question answering (Wang et al., 2023c; Xiong et al., 2024; Jeong et al., 2024). The most relevant works are Ye et al. (2021); Naik et al. (2022), which leverage clinical literature to augment clinical predictive models. Compared to these works, our contribution lies in two folds: (1) we design a retrieval augmentation for structured EHRs with a diverse collection of external knowledge, which provides more relevant information for target clinical prediction tasks; (2) we incorporate a co-training scheme to leverage both the visit-level information and external knowledge for predictions.
@@
Many studies attempt to harness external knowledge for clinical prediction tasks. The majority of them leverage structured knowledge, such as medical ontology (Choi et al., 2017; Panigutti et al., 2020; Lu et al., 2021), to capture hierarchical relationships among medical codes, or employ personalized knowledge graphs (Xu et al., 2023b; Jiang et al., 2024) to integrate patient-specific information. However, these methods often suffer from limited coverage of all medical codes due to the complexity of surface names. Alternatively, some approaches utilize unstructured medical text for health prediction tasks (Ye et al., 2021). However, Ye et al. (2021) rely on a restricted corpus of approximately 30,000 passages as their external corpus, resulting in limited coverage. Besides, it only considers ICD codes and uses the Levenshtein distance for retrieval, which cannot fully capture the semantics of medical codes.
@@ 3.1 Problem Setup
The EHR data consists of a group of patients $\mathcal{P}$ with corresponding hospital visits $V=\{v_{1},v_{2},...,v_{|V|}\}$. Each visit $v_{i}$ includes a set of medical codes $C_{i}\subset\mathcal{C}$, where $\mathcal{C}$ is the total set of medical codes for $\mathcal{P}$. In this study, $\mathcal{C}$ contains multiple types of medical codes including diseases, medications, and procedures. Each medical code $c_{i}\in C_{i}$ is a clinical concept, and it is associated with a name $s_{i}$ in the form of short text snippets. Given the clinical record $v_{i}$ with the involved medical codes $\mathcal{C}_{i}$, we aim to predict the patient’s clinical outcome $y_{i}$ (a binary label).
@@ 3.2 Retrieval Augmentation w/ Medical Codes
Multi-source Corpus Creation. Retrieval augmentation requires additional corpora as external knowledge. To ensure the coverage of clinical knowledge, we collect a diverse external resources =$\mathcal{M}=\{d_{1},d_{2},\ldots,d_{|\mathcal{M}|}\}$. We represent each knowledge unit as a raw text to facilitate retrieval. The detailed information of $\mathcal{M}$ is in Appendix B.
@@ 4.1 Experiment Setups
Datasets. We conduct experiments on the public MIMIC-III dataset (Johnson et al., 2016) and a private Ih-D1 dataset collected from a large healthcare system in the United States. We perform a 25-label phenotypes prediction task on MIMIC-III, and a cardiovascular disease (CVD) endpoints prediction task for diabetes patients on Ih-D1. We randomly split them into train/validation/test sets by 7:1:2. Please refer to Appendix A for details.
@@
  • HyGT (Cai et al., 2022): It leverages hypergraph transformers that regard patients as hyperedges and medical codes as nodes for EHR predictive tasks.
@@
  • SeqCare (Xu et al., 2023b): It is one of the strongest models for knowledge-graph-based EHR prediction with pruned graph structure learning and debiased training techniques.
@@
Implementation Details. In this work, we use Dragon (Lin et al., 2023) as the retriever for both passage $R_{D}(\cdot)$ and queries $R_{Q}(\cdot)$ [^1] https://huggingface.co/facebook/dragon-plus-{query,context}-encoder. We set $k=5$ during the retrieval stage. We choose UMLS-BERT (Michalopoulos et al., 2021) as the augmented model $g_{\phi}$ with 110M parameters for Ram-EHR and relevant baselines as $g_{\phi}$, with a maximum length of 512, and HyGT (Cai et al., 2022) as $f_{\theta}$ in main experiments. We also evaluate the effects of $g_{\phi}$ on PubmedBERT (Tinn et al., 2021), Clin-MobileBERT, and Clin-DistilBERT (Rohanian et al., 2023), as well as the effects of $f_{\theta}$ on GCT (Choi et al., 2020) and Transformer (Li et al., 2020) in Section 4.3. We set the learning rate to 5e-5 for $g_{\phi}$ and 1e-4 for $f_{\theta}$, batch size to 32, and the number of epochs to 5. We select $\beta,\lambda$ based on the performance of the validation set. All the experiments are conducted on a Linux server with one NVIDIA A100 GPU.
@@ 4.3.4 Parameter Study
In Figure 4, we conduct parameter studies on both datasets for $\beta$ and $\lambda$ in Eq. 5. Figure 4(a) demonstrates that the model achieves the best performance when $\beta$ is set to 0.2 and 0.4 on MIMIC-III and Ih-D1, respectively, while the gain diminishes at the extremes. This highlights the contribution of combining the predictions from both the augmented model and the local model on the performance gain. In addition, $\lambda$ is set to 1 and 5 on MIMIC-III and Ih-D1, respectively, according to Figure 4(b). The positive values of $\lambda$ indicate that the consistency loss enhances model performance.
@@ 4.4 Case Study
Figure 5(a) presents a case study on Ih-D1 to compare knowledge summarized by Ram-EHR and directly generated by LLM prompting. We observe that Ram-EHR provides more relevant information for the downstream task, particularly regarding the CVD outcome in this case, compared to direct LLM prompting. This also aligns with the human study evaluating the quality of 40 randomly sampled knowledge per type of code on a scale of [0,1,2] in Figure 5(b). Specifically, a score of 0 represents irrelevant knowledge, 1 denotes somewhat useful knowledge and 2 indicates highly useful knowledge. The detailed guidelines provided for the annotators are presented in Appendix E.
@@ Acknowledgement
One potential ethical consideration concerns the use of credential data (MIMIC-III and Ih-D1) with GPT-based online services. We have signed and strictly adhered to the PhysioNet Credentialed Data Use Agreement [^2] https://physionet.org/about/licenses/physionet-credentialed-health-data-license-150/ for the legal usage of the MIMIC-III dataset. To prevent sensitive information from being shared with third parties through APIs, we carefully follow the guidelines [^3] https://physionet.org/news/post/gpt-responsible-use for the responsible use of MIMIC data in online services. Specifically, we have requested to opt out of human review of the data by filling out the Azure OpenAI Additional Use Case Form [^4] https://aka.ms/oai/additionalusecase in order to utilize the Azure Open AI service while ensuring that Microsoft does not have access to the patient data. The utilization of LLMs in our framework is strictly to build medical concept-specific KGs. In addition, the building of medical concept-specific KGs does not involve direct interaction with any individual patient information. We iterate through all concepts in the medical coding system (e.g., CCS and ICD) to generate their respective KGs using LLMs, and these KGs are stored locally.
Appendix A Task Information
We present the detailed statistics of MIMIC-III and Ih-D1 in Table 3.
MIMIC-III.
@@ Appendix A Task Information
Ih-D1.
For the Ih-D1 dataset, we conduct a CVD outcome prediction task, which predicts whether patients with type 2 diabetes will experience CVD complications within 1 year after their initial diagnosis, including coronary heart disease (CHD), congestive heart failure (CHF), myocardial infarction (MI), or stroke. Diseases are identified by their ICD-9 or ICD-10 clinical codes. The usage of data has been approved by the Institutional Review Board (IRB).
Table 3: The statistics of MIMIC-III and Ih-D1.
Stats MIMIC-III Ih-D1
# of diagnosis 846 7915
# of medication 4525 489
# of procedure 2032 4321
# of health records 12353 36611
Appendix B Knowledge Sources
B.1 Descriptions
  • KG [^11] https://github.com/mims-harvard/PrimeKG (Chandak et al., 2023): We use PrimeKG in our experiments. It offers a comprehensive overview of diseases, medications, side effects, and proteins by merging 20 biomedical sources to detail 17,080 diseases across ten biological levels. For this study, we select knowledge triplets that contain medical codes within three types (disease, medication, procedure) used in this work, resulting in 707k triplets. We use the template in Appendix B.2 to transform these triplets into sentences.
B.2 Translating Format
@@ Appendix F Cost Information
Utilizing GPT-3.5-turbo as our base LLM model for generating summarized knowledge, we observe an average retrieval augmentation cost of \$0.0025 per medical code in MIMIC-III and \$0.0032 in Ih-D1. Consequently, Ram-EHR does not result in excessive monetary expenses.

### v3

@@ Abstract
Ran Xu1thanks:  Equal contribution., Wenqi Shi211footnotemark: 1, Yue Yu2, Yuchen Zhuang2, Bowen Jin3 May D. Wang2, Joyce C. Ho1, Carl Yang1 1 Emory University  2 Georgia Institute of Technology 3 University of Illinois at Urbana Champaign {ran.xu,joyce.c.ho,j.carlyang}@emory.edu, {wshi83,yueyu,yczhuang,maywang}@gatech.edu, bowenj4@illinois.edu
@@ 1 Introduction
To further improve the downstream predictive performance, several works attempt to augment the EHR visits with external knowledge. For example, van Aken et al. (2021) and Naik et al. (2022) incorporate additional clinical notes, although these clinical notes can be noisy and contain irrelevant contents for clinical predictions; another solution is to leverage external clinical knowledge graphs (KGs), such as UMLS (Chandak et al., 2023), which contains rich medical concepts (e.g., disease, medications) and their corresponding relationships (Jin et al., 2023). Integrating KGs with EHRs has been shown to boost model performance (Choi et al., 2017; Xu et al., 2023c; Gao et al., 2023). However, these works mostly rely on knowledge from a single source and medical KGs mainly focus on specific types of relations (e.g., hierarchical relations), which do not comprehensively capture the semantic information for medical codes (e.g., phenotype). Besides, it is non-trivial to align medical codes in EHRs with KGs due to the non-uniformity of surface names (e.g., abbreviations or colloquial terms) (Hao et al., 2021; Zhang et al., 2022). Jiang et al. (2024) use knowledge generated from large language models (LLMs) to assist EHR prediction, but LLMs may not always provide the most relevant knowledge for target tasks and face the risk of hallucination. Effectively leveraging external knowledge to facilitate EHR predictive tasks remains a significant challenge.
@@ 2 Related Works
Retrieval augmented learning, which collects additional contextual information from external corpus, has shown effectiveness on diverse tasks including language modeling (Borgeaud et al., 2022; Wang et al., 2023a), knowledge-intensive NLP (Lewis et al., 2020; Shi et al., 2023a), commonsense reasoning (Yu et al., 2022a), code genereation (Parvez et al., 2021), and few/zero-shot learning (Xu et al., 2023b; Izacard et al., 2023). Compared to the general domain, the application of retrieval-augmented learning to clinical tasks is still under-explored. Some efforts have been paid to retrieval augmented clinical language models (Zakka et al., 2024), clinical decision making (Shi et al., 2023b; Thompson et al., 2023) and clinical question answering (Wang et al., 2023c; Xiong et al., 2024; Jeong et al., 2024). The most relevant works are Ye et al. (2021); Naik et al. (2022), which leverage clinical literature to augment clinical predictive models. Compared to these works, our contribution lies in two folds: (1) we design a retrieval augmentation for structured EHRs with a diverse collection of external knowledge, which provides more relevant information for target clinical prediction tasks; (2) we incorporate a co-training scheme to leverage both the visit-level information and external knowledge for predictions.
@@
Many studies attempt to harness external knowledge for clinical prediction tasks. The majority of them leverage structured knowledge, such as medical ontology (Choi et al., 2017; Panigutti et al., 2020; Lu et al., 2021), to capture hierarchical relationships among medical codes, or employ personalized knowledge graphs (Xu et al., 2023c; Jiang et al., 2024) to integrate patient-specific information. However, these methods often suffer from limited coverage of all medical codes due to the complexity of surface names. Alternatively, some approaches utilize unstructured medical text for health prediction tasks (Ye et al., 2021). However, Ye et al. (2021) rely on a restricted corpus of approximately 30,000 passages as their external corpus, resulting in limited coverage. Besides, it only considers ICD codes and uses the Levenshtein distance for retrieval, which cannot fully capture the semantics of medical codes.
@@ 3.1 Problem Setup
The EHR data consists of a group of patients $\mathcal{P}$ with corresponding hospital visits $V=\{v_{1},v_{2},...,v_{|V|}\}$. Each visit $v_{i}$ includes a set of medical codes $C_{i}\subset\mathcal{C}$, where $\mathcal{C}$ is the total set of medical codes for $\mathcal{P}$. In this study, $\mathcal{C}$ contains multiple types of medical codes including diseases, medications, and procedures. Each medical code $c_{i}\in C_{i}$ is a clinical concept, and it is associated with a name $s_{i}$ in the form of short text snippets. Given the clinical record $v_{i}$ with the involved medical codes $C_{i}$, we aim to predict the patient’s clinical outcome $y_{i}$ (a binary label).
@@ 3.2 Retrieval Augmentation w/ Medical Codes
Multi-source Corpus Creation. Retrieval augmentation requires additional corpora as external knowledge. To ensure the coverage of clinical knowledge, we collect a diverse external resources =$\mathcal{M}=\{d_{1},d_{2},\ldots,d_{|\mathcal{M}|}\}$. We represent each knowledge unit as a raw text to facilitate retrieval, described as follows:
  • KG [^6] https://github.com/mims-harvard/PrimeKG (Chandak et al., 2023): We use PrimeKG in our experiments. It offers a comprehensive overview of diseases, medications, side effects, and proteins by merging 20 biomedical sources to detail 17,080 diseases across ten biological levels. For this study, we select knowledge triplets that contain medical codes within three types (disease, medication, procedure) used in this work, resulting in 707k triplets. We use the template in Appendix B to transform these triplets into sentences.
@@ 4.1 Experiment Setups
Datasets. We present the detailed statistics of MIMIC-III and Cradle in Table 1.
MIMIC-III.
Cradle.
For the Cradle dataset, we conduct a CVD outcome prediction task, which predicts whether patients with type 2 diabetes will experience CVD complications within 1 year after their initial diagnosis, including coronary heart disease (CHD), congestive heart failure (CHF), myocardial infarction (MI), or stroke. Diseases are identified by their ICD-9 or ICD-10 clinical codes. The usage of data has been approved by the Institutional Review Board (IRB).
Table 1: The statistics of MIMIC-III and Cradle.
Stats MIMIC-III Cradle
# of diagnosis 846 7915
# of medication 4525 489
# of procedure 2032 4321
# of health records 12353 36611
@@
  • HyGT (Xu et al., 2023a): It leverages hypergraph transformers that regard patients as hyperedges and medical codes as nodes for EHR predictive tasks.
@@
  • SeqCare (Xu et al., 2023c): It is one of the strongest models for knowledge-graph-based EHR prediction with pruned graph structure learning and debiased training techniques.
@@
Implementation Details. In this work, we use Dragon (Lin et al., 2023) as the retriever for both passage $R_{D}(\cdot)$ and queries $R_{Q}(\cdot)$ [^8] https://huggingface.co/facebook/dragon-plus-{query,context}-encoder. We set $k=5$ during the retrieval stage. We choose UMLS-BERT (Michalopoulos et al., 2021) as the augmented model $g_{\phi}$ with 110M parameters for Ram-EHR and relevant baselines as $g_{\phi}$, with a maximum length of 512, and HyGT (Xu et al., 2023a) as $f_{\theta}$ in main experiments. We also evaluate the effects of $g_{\phi}$ on PubmedBERT (Tinn et al., 2021), Clin-MobileBERT, and Clin-DistilBERT (Rohanian et al., 2023), as well as the effects of $f_{\theta}$ on GCT (Choi et al., 2020) and Transformer (Li et al., 2020) in Section 4.3. We set the learning rate to 5e-5 for $g_{\phi}$ and 1e-4 for $f_{\theta}$, batch size to 32, and the number of epochs to 5. We select $\beta,\lambda$ based on the performance of the validation set. All the experiments are conducted on a Linux server with one NVIDIA A100 GPU.
@@ 4.3.4 Parameter Study
In Figure 4, we conduct parameter studies on both datasets for $\beta$ and $\lambda$ in Eq. 5. Figure 4(a) demonstrates that the model achieves the best performance when $\beta$ is set to 0.2 and 0.4 on MIMIC-III and Cradle, respectively, while the gain diminishes at the extremes. This highlights the contribution of combining the predictions from both the augmented model and the local model on the performance gain. In addition, $\lambda$ is set to 1 and 5 on MIMIC-III and Cradle, respectively, according to Figure 4(b). The positive values of $\lambda$ indicate that the consistency loss enhances model performance.
@@ 4.4 Case Study
Figure 5(a) presents a case study on Cradle to compare knowledge summarized by Ram-EHR and directly generated by LLM prompting. We observe that Ram-EHR provides more relevant information for the downstream task, particularly regarding the CVD outcome in this case, compared to direct LLM prompting. This also aligns with the human study evaluating the quality of 40 randomly sampled knowledge per type of code on a scale of [0,1,2] in Figure 5(b). Specifically, a score of 0 represents irrelevant knowledge, 1 denotes somewhat useful knowledge and 2 indicates highly useful knowledge. The detailed guidelines provided for the annotators are presented in Appendix E.
@@ Acknowledgement
One potential ethical consideration concerns the use of credential data (MIMIC-III and Cradle) with GPT-based online services. We have signed and strictly adhered to the PhysioNet Credentialed Data Use Agreement [^9] https://physionet.org/about/licenses/physionet-credentialed-health-data-license-150/ for the legal usage of the MIMIC-III dataset. To prevent sensitive information from being shared with third parties through APIs, we carefully follow the guidelines [^10] https://physionet.org/news/post/gpt-responsible-use for the responsible use of MIMIC data in online services. Specifically, we have requested to opt out of human review of the data by filling out the Azure OpenAI Additional Use Case Form [^11] https://aka.ms/oai/additionalusecase in order to utilize the Azure Open AI service while ensuring that Microsoft does not have access to the patient data. The utilization of LLMs in our framework is strictly to build medical concept-specific KGs. In addition, the building of medical concept-specific KGs does not involve direct interaction with any individual patient information. We iterate through all concepts in the medical coding system (e.g., CCS and ICD) to generate their respective KGs using LLMs, and these KGs are stored locally.
Appendix A Phenotypes
@@
The details of 25 phenotypes are listed in table 3.
Appendix B Knowledge Translating Format
@@ Appendix F Cost Information
Utilizing GPT-3.5-turbo as our base LLM model for generating summarized knowledge, we observe an average retrieval augmentation cost of \$0.0025 per medical code in MIMIC-III and \$0.0032 in Cradle. Consequently, Ram-EHR does not result in excessive monetary expenses.